OpenStudy (anonymous):

does anyone know how to solve this; ax + 2xy = 14 for y

jimthompson5910 (jim_thompson5910):

Your first step is to subtract ax from both sides

jimthompson5910 (jim_thompson5910):

What does that give you?

OpenStudy (anonymous):

but what happens if u subtract 14-ax?? what would that give u?

jimthompson5910 (jim_thompson5910):

you're just subtracting ax from both sides , not 14-ax

OpenStudy (anonymous):

okay then forget about it.. ima just go to tutoring tommorow lol

jimthompson5910 (jim_thompson5910):

alright, sounds like a good idea

OpenStudy (anonymous):

lol yup it is >.<

OpenStudy (anonymous):

ax + 2xy = 14 for y -ax -ax 2xy=14-ax /2x /2x y=(14-ax)/2x